A physical-fitness association in Calgary Alberta is including the mile run in its secondary-school fitness test for girls. The time for this event is approximately normally distributed with a mean of 400 seconds and a standard deviation of 40 seconds. If the association wants to select the fastest 10% as “good”, what is the maximum time (in seconds) a runner will need in order to qualify as “good”? Select one: a. 334.2 b. 501.2 c. 400 d. 348.8 e. 451.2
Answer #1

solution:

Given that,

mean = = 400

standard deviation = =40

Using standard normal table,

P(Z > z) = 10%

= 1 - P(Z < z) = 0.10

= P(Z < z ) = 1 - 0.10

= P(Z < z ) = 0.90

= P(Z < z ) = 0.90  

z = 1.28 (using standard normal (Z) table )

Using z-score formula  

x = z * +

x= 1.28 *40+400

x= 451.2